OpenStudy (anonymous):

Which is equal to 6sqrt32^3 By 6 sqrt I mean the root is 6, not 6 * sqrt 32^2. I'm in logarithms and I checked the lesson.. it's not there D:

OpenStudy (anonymous):

Answer choices: \[4\sqrt{2}\] \[4\sqrt{3}\] \[16\sqrt{2}\] \[16\sqrt{3}\]

OpenStudy (agl202):

6√(48^3) = 1152√3 But this question is so far off the mark I'm guessing you got the terminology wrong, and you meant to ask which is equal to the sixth root of 48^3. If that's the case, then (48^3)^(1/6) = [(2^4 * 3)^3]^(1/6) = (2^12 * 3^3)^(1/6) = 2^2 √3 = 4√3 B.

OpenStudy (anonymous):

But that's the answer to 48^3. I need 32^3.

OpenStudy (agl202):

Oh... k dunno

OpenStudy (anonymous):

It's fine :)

OpenStudy (agl202):

Sorry :(

OpenStudy (anonymous):

Really it's ok, I'll get the answer somehow. :)